# Semi-orthogonality in Fukaya-Seidel mirrors to blowups of abelian varieties

### v1

@@ 4. Product structure
$\operatorname{Ext}(\pi^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}}),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}}))\otimes\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}}))$
$\hskip 72.26999pt\to\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}})).$
These can be simplified using the isomorphism of Equation (2.3)
$\operatorname{Ext}(p^{*}i^{*}(\mathcal{L}^{k}\boxtimes\mathcal{O}_{\mathbb{C}}),j^{*}\pi^{*}(\mathcal{L}^{j}\boxtimes\mathcal{O}_{\mathbb{C}})[-1])\cong\operatorname{Ext}(\mathcal{L}^{k}|_{H},\mathcal{L}^{j}|_{H})[-1]$
@@
under which we restrict to $H\times\{0\}$. For example, sections of $\mathcal{O}_{\mathbb{C}}$ are polynomials, and under this isomorphism they are evaluated at 0. Sections of $\mathcal{L}^{j}$ for $j>0$ are multi-theta functions, infinite series, as in [ACLL, Equation (2.14)].
When composing morphisms between factors on the A-side, of the three Lagrangians, we see similar behavior; by the geometry of wrapping we can have either two which are arcs and one which is U-shaped, or two which are U-shaped and one which is an arc. For this we would like to use the result of [ACLL23b, Corollary 2.6], therefore we need to prove the assumption in the corollary, that given a pseudo-holomorphic curve $u$ counted in the composition map over a triangle away from the singularity, and given the map $u^{\prime}$ defined in [ACLL23a, Proposition 6.10] obtained by moving one Lagrangian under a Lagrangian isotopy so $\partial u^{\prime}$ is in a fiber, then $u^{\prime}]$ is relatively homologous to a pseudo-holomorphic triangle.
4.1. Grading
Let $L$ be $L_{j_{2}}$ (respectively $K_{j_{2}}$). For $\mathcal{D}(Y^{0},(L_{j_{1}},L,K_{j_{3}}),(p_{0},p_{1},p_{2}))$ the class of discs mapping to the Lagrangian boundary conditions (see [ACLL23a, Definition 5.1]), let
$\mathcal{M}(Y,(L_{j_{1}},L,K_{j_{3}}),(p_{0},p_{1},p_{2});\beta,J)\\ =\{u\in\mathcal{D}(Y^{0},(L_{j_{1}},L,K_{j_{3}}),(p_{0},p_{1},p_{2}))\mid[u]=\beta,\overline{\partial}_{J}(u)=0\}$
@@ 4.1. Grading
be the moduli space of $J$-holomorphic discs in the composition
(4.1) $HF(L,K_{j_{3}}[-1])\otimes HF(L_{j_{1}},L)\to HF(L_{j_{1}},K_{j_{3}}[-1])$
depicted under projection of $W_{0}$ in Figure 7 when $L=L_{j_{2}}$ (respectively in Figure 9 when $L=K_{j_{2}}$).
The typical grading in a bigon in $\mathbb{R}^{2}$ has degree 0 on the right intersection point and degree 1 on the left intersection point, as in Figure 5 and [Aur14, Example 1.9]. When grading the intersection points in a Floer complex $CF(L_{1},L_{2})$ in the total space of a Landau-Ginzburg model, the fiber and base contributions add by [ACLL23b, Lemma 3.11 and Corollary 3.12]. The fiber case is covered in [ACLL23a, Example 4.6], where $\ell_{k}$ is linear of slope $k$ and the Lagrangian grading is $\frac{1}{\pi}\arctan(k)$. So here we focus on the base. We make a continuous choice of angle $\alpha_{j}(t)$ along the projected Lagrangian $W_{0}(L_{j})$ for a tangent vector to $T_{\gamma_{L_{j}(t)}}W_{0}(L_{j})$ relative to the positive horizontal line field, indicated in Figures 7 and 9 by arrows.
To find the degree of an intersection point in the base, we can break it up into three steps. Step 1 is to find $\alpha_{2}(t)-\alpha_{1}(t)$ at values of $t$ where the paths intersect (direction of the arrows on the base Lagrangians matter). Step 2: at that same value of $t$, we rotate the line $T_{\gamma_{L_{2}(t)}}W_{0}(L_{2})$ counterclockwise until it coincides with that of $T_{\gamma_{L_{1}(t)}}W_{0}(L_{1})$ (direction of the arrows on the base Lagrangians don’t matter). Since counterclockwise is always positive, we add that to $\alpha_{2}(t)-\alpha_{1}(t)$. This is referred to as the reverse canonical short path. Step 3: Twice this winding number is the degree, because rotating a line back to itself, through an angle $\pi$, should count as one rotation.
Remark 4.1.
Note that $(\exp(2\pi i\alpha_{j}(t)))^{2}=\alpha_{L}^{\text{Hor}}(p)$ of [ACLL23b, Lemma 3.11, Eq. (3.8)] where, for horizontal lift $(\dot{\gamma}_{L_{j}}(t))^{\#}$ and recalling $r(t)$ and $\theta(t)$ parametrizing $L_{j}$ as in Definition 3.4,
For example for a circle $\dot{r}(t)=0$, $\dot{\theta}(t)=1$ and
implies one choice for $\alpha_{j}(t)$ is
$\alpha_{j}(t)=\theta(t)+1/2.$
In contrast to the typical grading, here we want choices so that the three bigons involved in each differentials in the three $HF$ factors appearing in the composition in Equation (4.1) have degree 0 on the left intersection point and degree 1-1 on the right intersection point. This is done in order to mirror the shift in Ext groups on the B-side. That is, U-shapes in the base are all graded the same which means in the bigon, the choice of $\alpha_{j_{2}}(t)$ is shifted by $\pi$ from Figure 5, as in Figure 7. Replace the choice of $\alpha_{2}(t)$ in Figure 5 with $\alpha_{j_{2}}(t)=\alpha_{2}(t)-\pi$ for the definition of $L_{j_{2}}$. We do a similar adjustment for $K_{j_{3}}$ so that $\alpha_{j_{3}}(t)=-\pi$. Thus $K_{j_{3}}$ is replaced with $K_{j_{3}}[-1]$.
Labeling the base intersection points as $b_{0},\ldots,b_{5}$, we calculate the degrees $\deg(b_{j})$. Explanation is given in Figure 8 for $b_{0}$ and $b_{1}$. The other points are calculated similarly.
[…]
Figure 5.
$L_{j_{1}}$
$L_{j_{2}}$
$K_{j_{3}}$
@@
Figure 6. Composition $HF(L_{j_{2}},K_{j_{3}}[-1])\otimes HF(L_{j_{1}},L_{j_{2}})\to HF(L_{j_{1}},K_{j_{3}}[-1])$ depicted in the base in the cylinder model. The thicker arrow corresponds to the grading on the input projected Lagrangian of the Floer group.
Figure 7.
[…]
Figure 8. : Add $\frac{1}{2}$ because the oriented tangent vector rotates counterclockwise $\frac{\pi}{2}$ from $b_{0}$ to $b_{1}$
@@
Figure 9. Composition $HF(K_{j_{2}}[-1],K_{j_{3}}[-1])\otimes HF(L_{j_{1}},K_{j_{2}}[-1])\to HF(L_{j_{1}},K_{j_{3}}[-1])$ depicted in the $\mathbb{C}^{*}$ base model, denote $D$ as the shaded region
@@
In the moduli space defining
$HF(L_{j_{2}},K_{j_{3}}[-1])\otimes HF(L_{j_{1}},L_{j_{2}})\to HF(L_{j_{1}},K_{j_{3}}[-1]),$
consider
(4.3) $u\in\mathcal{M}(Y,(L_{j_{1}},L_{j_{2}},\Phi_{H}(K_{j_{3}}[-1])),(p_{0},p_{1},p_{2});\beta,J)$
where $\Phi_{H}:Y^{0}\to Y^{0}$ is a Hamiltonian diffeomorphism to be defined in the proof and $W_{0}(\beta)$ covers the shaded region in Figure 7. Let $u^{\prime}$ be obtained by Lagrangian isotopy of $K_{j_{3}}$ so that all three intersection points pass through the same fiber, following [ACLL23b, Corollary 2.6]. Then $u^{\prime}]$ is relatively homologous to a pseudo-holomorphic triangle $u^{\prime\prime}$ in the fiber bounded by $\ell_{j_{1}}\cup\ell_{j_{2}}\cup\ell_{j_{3}}$.
The point is, we want to get a handle on $u$ to consider composition on the A-side, but there is ambiguity in how we draw the Lagrangians. With this Lemma and then subsequently [ACLL23b, Corollary 2.6], we can move everything to a fiber triangle $u^{\prime\prime}$, and then count the triangles by area from HMS on tori [ACLL, §4.3], [Fuk02].
@@
We construct a singular 3-chain, whose boundary is $u^{\prime}]-[u^{\prime\prime}]$ where $u^{\prime\prime}$ is in a fiber. The 3-chain is a continuous map from the unit cube to $Y^{0}$, and a cube can be triangulated into 3-simplices. For simplicity, denote $W_{0}$ as $v$.
Let $\delta(s,t):I\times I\to\mathbb{C}$ for $I=[0,1]$ parametrize the projection of the triangle $(v\circ u)(\mathbb{D})$ so that $\delta:(0,1)\times(0,1)\to(v\circ u)(\text{int}(\mathbb{D}))$ is a homeomorphism, $\delta(1,t),0\leq t\leq 1$ is the top side of the base triangle in Figure 7 before the isotopy, and the isotopy happens over the curve $\delta(s,t),0\leq s\leq 1$ for each $t$ (this is consistent with the use of $\delta$ in [ACLL23b, Corollary 2.6]). Then for $\Phi_{\gamma(0)\to\gamma(1)}$ denoting parallel transport along a path $\gamma$ between the fibers over $\gamma(0)\to\gamma(1)$ (as in [ACLL23b, Equation (1.3)]), define the singular 3-chain
(4.4) $Q:I\times I\times I$ $\to Y$
@@
where
$D:=(v\circ u)(\mathbb{D})$
and $u_{D}:D\to Y^{0}$ is defined by
$u_{D}(\delta(s,t))=u(z(s,t))$
where $z(s,t)\in\mathbb{D}$ is uniquely defined by
$(v\circ u)(z(s,t))=\delta(s,t).$
because $u$ is 1-1. This is because we are working with a Landau-Ginzburg model [Sei08, Equation (17.2)] where curves in structure maps are sections of the superpotential $v$. In other words, $u_{D}:D\to Y^{0}$ is a $((v\circ u)_{*}j,J)$-holomorphic section of $v$ if $u$ was a $(j,J)$-holomorphic curve.
Then $Q$ is a continuous map by definition. The image of $Q$ is depicted in [ACLL23b, Figure 2], the prism over the middle image. Its boundary is (where paranthetical comments refer to [ACLL23b, Figure 2])
(4.5) $z=0,\quad 0\leq s,t\leq 1$ $\mapsto[u]\text{ (bottom triangle) }$
$z=1,\quad 0\leq s,t\leq 1$ $\mapsto[u^{\prime\prime}]\text{ (fiber triangle) }$
$s=0,\quad 0\leq z,t\leq 1$ $\mapsto[u^{\prime}]-[u]\text{ (top quadrilateral obtained from the Lagrangian isotopy) }$
$s=1,\quad 0\leq z,t\leq 1$ $\mapsto u(\delta(1,-))\text{ (constant at bottom vertex) }$
$t=0,1,\quad 0\leq s,z\leq 1$ $\mapsto K_{j_{2}}\cup K_{j_{3}}\text{ (front and back triangles) }$
where $u^{\prime\prime}$ is defined as follows. In the case of Figure 7, define the parameters along the boundary of the triangle in the base by
(4.6) $\gamma_{L_{j_{1}}}(t_{L_{j_{1}}}(s))$ $=\delta(s,1)$
$\gamma_{L_{j_{2}}}(t_{L_{j_{2}}}(s))$ $=\delta(s,0)$
$\gamma_{K_{j_{3}}}(t_{K_{j_{3}}}(t))$ $=\delta(0,t).$
This clarifies the relation between the homotopy $\delta(s,t)$ and the parameters defining the base curves for the Lagrangians. Then the $z=1$ face has boundary on one side on the Lagrangian in the fiber over S-S given by
@@
For the third side, we apply a Hamiltonian diffeomorphism $\Phi_{H}$ defined first on fibers over the initial third side
(4.9) $\Phi_{H}=\Phi_{\delta(1,t)\to\delta(0,t)}\circ\Phi_{\delta_{K_{j_{3}}}(0)\to\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}^{-1}:\bigcup_{0\leq t\leq 1}Y^{0}_{\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}$ $\mapsto\bigcup_{0\leq t\leq 1}Y^{0}_{\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}$
$\Phi_{\delta_{K_{j_{3}}}(0)\to\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}(p)$ $\mapsto\Phi_{\delta(1,t)\to\delta(0,t)}(p).$
Because the composition of the two maps is parallel transport around a closed loop not enclosing the singular fiber by Equation (4.6), the composition is Hamiltonian (for the submanifold given by the fibers over the $\delta(0,t)$ part of the triangle), by [MS98, Theorem 6.21 (ii)]. This defines a Hamiltonian function $H$ on that submanifold, which we can then extend to the whole space smoothly. We call the resulting Hamiltonian diffeomorphism $\Phi_{H}$. The reason for applying this Hamiltonian diffeomorphism is so that under the $\delta$ isotopy, the third side is in the S-S fiber again in a linear Lagrangian

### v2

@@ 3.5. Product structure
$\text{Hom}_{D^{b}(X)}(D^{b}(V\times\mathbb{C}),D^{b}(V\times\mathbb{C}))\otimes\text{Hom}_{D^{b}(X)}(D^{b}(H),D^{b}(V\times\mathbb{C}))\to\text{Hom}_{D^{b}(X)}(D^{b}(H),D^{b}(V\times\mathbb{C})).$
Other compositions involving both $H$ and $V\times\mathbb{C}$ factors are 0. More specifically, we consider
(3.65) $\operatorname{Ext}^{d_{2}}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}}))$
$\otimes\operatorname{Ext}^{d_{1}}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)))$
$\hskip 72.26999pt\to\operatorname{Ext}^{d_{1}+d_{2}}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}}))$
which reduces to
(3.66) $\operatorname{Ext}^{d_{2}}(\mathcal{L}^{k_{2}}|_{H},\mathcal{L}^{k_{3}}|_{H})[-1]\otimes\operatorname{Ext}^{d_{1}}(\mathcal{L}^{k_{1}}|_{H},\mathcal{L}^{k_{2}}|_{H})\to\operatorname{Ext}^{d_{1}+d_{2}}(\mathcal{L}^{k_{1}}|_{H},\mathcal{L}^{k_{3}}|_{H})[-1]$
and
$\operatorname{Ext}^{d_{2}}(\pi^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}}),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}}))\otimes\operatorname{Ext}^{d_{1}}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}}))$
(3.67) $\hskip 72.26999pt\to\operatorname{Ext}^{d_{1}+d_{2}}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}}))$
which reduces to
(3.68) $\operatorname{Ext}^{d_{2}}(\mathcal{L}^{k_{2}}|_{H},\mathcal{L}^{k_{3}}|_{H})\otimes\operatorname{Ext}^{d_{1}}(\mathcal{L}^{k_{1}}|_{H},\mathcal{L}^{k_{2}}|_{H})[-1]\to\operatorname{Ext}^{d_{1}+d_{2}}(\mathcal{L}^{k_{1}}|_{H},\mathcal{L}^{k_{3}}|_{H})[-1]$
where we used the isomorphisms of Equations (2.2) and (2.3)
$\operatorname{Ext}_{X}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}}))\cong\operatorname{Ext}_{H}(\mathcal{L}^{k_{1}}|_{H},\mathcal{L}^{k_{2}}|_{H})[-1]$
@@
under which we restrict to $H\times\{0\}$. Sections of $\mathcal{L}^{j}$ for $j>0$ are multi-theta functions, infinite series, as in [ACLL_tori, Equation (2.14), $g=2$]. These are restricted to $H$. Sections of $\mathcal{O}_{\mathbb{C}}$, polynomials, are evaluated at 0. We calculate $\text{Ext}^{*}(\mathcal{L}^{i}|_{H},\mathcal{L}^{j}|_{H})\cong H^{*}(\mathcal{L}^{j-i}|_{H})$ by using the long exact sequence associated to tensoring the short exact sequence
(3.69) $0\to\mathcal{L}^{-1}\to\mathcal{O}_{V}\to\iota_{*}\mathcal{O}_{H}\to 0$
with $\mathcal{L}^{\otimes(j-i)}$, for $\iota:H\to V$. Taking l=jil=j-i and for brevity letting $\mathcal{L}^{\otimes l}$ be denoted as $\mathcal{L}^{l}$, we find that the cohomology groups of $\mathcal{L}^{l}|_{H}$ can be expressed in terms of cohomology of line bundles on $V$ in [ACLL_tori, Theorem 3.3]. We have
(3.70) $H^{d}(\mathcal{L}^{l}|_{H})\cong\begin{cases}\delta_{d0}\cdot\text{coker}(H^{0}(\mathcal{L}^{l-1})\xrightarrow[]{\otimes\vartheta}H^{0}(\mathcal{L}^{l}))&l>1\\ H^{d+1}(\mathcal{O}_{V})&l=1\\ H^{d}(\mathcal{O}_{V})&l=0\\ \delta_{d1}\cdot\text{ker}(H^{1}(\mathcal{L}^{l-1})\xrightarrow[]{\otimes\vartheta}H^{1}(\mathcal{L}^{l}))&l<0\end{cases},\qquad 0\leq d\leq 1$
Composition is given by multiplication of theta functions [ACLL_tori, §3.3].
A-side. Composition in the Fukaya category is defined by counting triangles weighted by area, see [ACLL_Bfield, Definition 5.13, $k=2$].
When composing two morphisms on the A-side, either the first morphism is on the $\mathcal{A}_{L}$ factor and then we compose with a morphism in $\text{Hom}(\mathcal{A}_{L},\mathcal{A}_{K})$, or the first morphism is in $\text{Hom}(\mathcal{A}_{L},\mathcal{A}_{K})$ and we compose with one in $\mathcal{A}_{K}$. Other compositions, involving both factors, are zero by semi-orthogonality. In other words, the three Lagrangians involved in a given composition will either have two which cover cotangent fiber arcs and one which covers a U-shape, or two which cover U-shapes and one which covers a cotangent fiber arc. Let $\mathbf{L}_{j}$ denote $L_{j}$ (respectively $K_{j}[1]$). For the moduli space of discs $\mathcal{D}(Y^{0},(L_{j_{1}},\mathbf{L}_{j_{2}},K_{j_{3}}[1]),(p_{0},p_{1},p_{2}))$ with Lagrangian boundary conditions defined in [ACLL_Bfield, Definition 5.1], let
(3.71) $\mathcal{M}(Y^{0},(L_{j_{1}},\mathbf{L}_{j_{2}},K_{j_{3}}[1]),(p_{0},p_{1},p_{2});\beta,J)=$
$\hskip 72.26999pt\{u\in\mathcal{D}(Y^{0},(L_{j_{1}},\mathbf{L}_{j_{2}},K_{j_{3}}[1]),(p_{0},p_{1},p_{2}))\mid[u]=\beta,\overline{\partial}_{J}(u)=0,E(u)<\infty\}$
@@
be the moduli space of $J$-holomorphic discs of class $\beta$ counted in the composition
(3.72) $HF^{*}(\mathbf{L}_{j_{2}},K_{j_{3}}[1])\otimes HF^{*}(L_{j_{1}},\mathbf{L}_{j_{2}})\to HF^{*}(L_{j_{1}},K_{j_{3}}[1])$
depicted under projection of $W_{0}$ in Figure 8 when $\mathbf{L}_{j_{2}}=L_{j_{2}}$ (respectively in Figure 9 when $\mathbf{L}_{j_{2}}=K_{j_{2}}[1]$).
$W_{0}(L_{j_{1}})$
$W_{0}(L_{j_{2}})$
$W_{0}(K_{j_{3}}[1])$
@@
o
Figure 8. Composition $HF(L_{j_{2}},K_{j_{3}}[1])\otimes HF(L_{j_{1}},L_{j_{2}})\to HF(L_{j_{1}},K_{j_{3}}[1])$ depicted in the base. The thicker arrow corresponds to the grading on the input projected Lagrangian of the Floer group.
@@
$W_{0}(L_{j_{1}})$
$W_{0}(K_{j_{2}}[1])$
$W_{0}(K_{j_{3}}[1])$
Figure 9. Composition $HF(K_{j_{2}}[1],K_{j_{3}}[1])\otimes HF(L_{j_{1}},K_{j_{2}}[1])\to HF(L_{j_{1}},K_{j_{3}}[1])$ depicted in the base.
Then we claim that these compositions can be calculated by moving one Lagrangian so that all three points in the base coincide, see Lemma 3.30. Then we can count triangles in the fiber, see Lemma 3.31.
@@
Consider a pseudo-holomorphic triangle $u\in\mathcal{M}(Y^{0},(L_{j_{1}},\mathbf{L}_{j_{2}},K_{j_{3}}[1]),(p_{0},p_{1},p_{2});\beta,J)$ counted in the composition
$HF^{*}(\mathbf{L}_{j_{2}},K_{j_{3}}[1])\otimes HF^{*}(L_{j_{1}},\mathbf{L}_{j_{2}})\to HF^{*}(L_{j_{1}},K_{j_{3}}[1]),$
where $W_{0}(\beta)$ covers the shaded region in Figure 8 for $\mathbf{L}_{j_{2}}=L_{j_{2}}$, respectively Figure 9 for $\mathbf{L}_{j_{2}}=K_{j_{2}}[1]$. Let $u^{\prime}$ be the new piecewise curve obtained after Lagrangian isotopy of $K_{j_{3}}[1]$, respectively $L_{j_{1}}$, so that all three intersection points pass through the same fiber.
Then
  1. (i)
    the image of $u^{\prime}$ is relatively homologous to a triangle $u^{\prime\prime}$ in a fiber, and
  2. (ii)
    after applying a Hamiltonian diffeomorphism to the isotoped Lagrangian, respectively also a symplectomorphism, $u^{\prime\prime}$ is bounded by $\ell_{j_{1}}\cup\ell_{j_{2}}\cup\ell_{j_{3}}$ in the fiber over S-S.
@@
(i) First assume $\mathbf{L}_{j_{2}}=L_{j_{2}}$. We will prove the statement by constructing a contractible singular 3-chain $Q$ whose boundary relative the Lagrangians is $u^{\prime}]-[u^{\prime\prime}]$. First we describe the construction in the base of $W_{0}$.
Let $\delta(s,t):I\times I\to\mathbb{C}$ for $I=[0,1]$ parametrize the projection of the triangle $W_{0}(u(\mathbb{D}))$, as follows. In the shaded triangle in Figure 8, on the boundary, the top of the square $I\times\{1\}$ maps to right side of the triangle along $W_{0}(L_{j_{2}})$, the left side of the square $\{0\}\times I$ maps to top of the triangle along $W_{0}(K_{j_{3}}[1])$, the bottom side of the square $I\times\{0\}$ maps to the the left of the triangle along $W_{0}(L_{j_{1}})$, and the right side of the square $\{1\}\times I$ maps to $b_{0}$. On its boundary, reparametrize the base curves defining the Lagrangians to match $\delta(s,t)$ along the boundary, so that
(3.73) $\delta(s,1)$ $=\gamma_{L_{j_{2}}}(t_{L_{j_{2}}}(s))$
$\delta(0,t)$ $=\gamma_{K_{j_{3}}}(t_{K_{j_{3}}}(t))$
$\delta(s,0)$ $=\gamma_{L_{j_{1}}}(t_{L_{j_{1}}}(s))$
$\delta(1,t)$ $=b_{0}.$
Extend to the interior so that $\delta:(0,1)\times(0,1)\to(W_{0}\circ u)(\text{int}(\mathbb{D}))$ is a homeomorphism. Then the Lagrangian isotopy of $K_{j_{3}}[1]$, in the base, contracts the curve $\{\delta(s,t)\}_{0\leq s\leq 1}$ for each fixed $t$. That is, at $s=0$, $\delta(0,t)\in W_{0}(K_{j_{3}}[1])$, while at $s=1$, $\delta(1,t)=b_{0}$.
Now we define the singular chain $Q$ in the total space as a continuous map from the unit cube to $Y^{0}$. Again we can replace $u$ with $u_{D}$ as in Equation (3.33), over the region in the base defined by $A(s,t)=\delta(s,t)$ here. We extend $\delta(s,t)$ to the total space by parallel transport. Recall $\Phi_{\gamma(0)\to\gamma(1)}$ denotes parallel transport along a path $\gamma$ between the fibers over $\gamma(0)\to\gamma(1)$, Definition 3.1. Define the singular 3-chain $Q$ to be $u_{D}$ on the bottom face $I\times I\times\{0\}$, the continuous map
(3.74) $u^{\prime\prime}=\Phi_{\delta(s,t)\to\delta(1,t)}\circ u_{D}$
to the fiber over $b_{0}$ on the top face $I\times I\times\{1\}$, and interpolate in between:
(3.75) $Q:I\times I\times I$ $\to Y^{0}$
@@
Then $Q$ is a continuous map by definition. The image of $Q$ is depicted in [ACLL_LG, Figure 2], the prism over the middle image. Note that the construction there is an example of $P$ in [ACLL_Bfield, Equation (6.11)] for LG models and describes the boundary of $Q$ constructed here. The map $u^{\prime}$ is defined by keeping $u$ and then “stretching” it further as we move the Lagrangian [ACLL_Bfield, Equation (6.11)]. On $\partial Q$, (paranthetical comments refer to [ACLL_LG, Figure 2])
(3.76) $h=0,\quad 0\leq s,t\leq 1$ $\mapsto[u]\text{ (bottom triangle) }$
$h=1,\quad 0\leq s,t\leq 1$ $\mapsto[u^{\prime\prime}]\text{ (fiber triangle) }$
$s=0,\quad 0\leq h,t\leq 1$ $\mapsto[u^{\prime}]-[u]\text{ (top quadrilateral obtained from the Lagrangian isotopy) }$
$s=1,\quad 0\leq h,t\leq 1$ $\mapsto u_{D}(b_{0})\text{ (constant at bottom vertex) }$
[…]
(ii) By Equation (3.74), the homology class $u^{\prime\prime}]\in H_{2}(T^{4})$ is in the fiber over S-S with boundary on the following Lagrangians:
@@
For the third Lagrangian, consider
(3.79) $\phi_{H_{\Delta}}=\Phi_{\delta(1,t)\to\delta(0,t)}\circ\Phi_{\delta_{K_{j_{3}}}(0)\to\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}^{-1}:\bigcup_{0\leq t\leq 1}Y^{0}_{\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}$ $\mapsto\bigcup_{0\leq t\leq 1}Y^{0}_{\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}$
$\Phi_{\delta_{K_{j_{3}}}(0)\to\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}(p)$ $\mapsto\Phi_{\delta(1,t)\to\delta(0,t)}(p),\;\;p\in Y^{0}_{-S}$
because recall that $\delta_{K_{j_{3}}}(0)=-S=b_{0}=\delta(1,t)$ by Definition 3.5. The two parallel transport maps in $\phi_{H_{\Delta}}$ compose to parallel transport around a closed loop, $\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))$ to $\delta_{K_{j_{3}}}(0)=-S$ backwards along $\delta_{K_{j_{3}}}$ and then $\delta(1,t)=-S$ backwards along the homotopy back to $\delta(0,t)$, which equals $\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))$ by Equation (3.73). As this loop does not enclose the singular fiber, the composition is Hamiltonian on $\bigcup_{0\leq t\leq 1}Y^{0}_{\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}$, by [symp_intro, Theorem 6.21(ii)], and corresponds to some Hamiltonian function which we call $H_{\Delta}$ to indicate it’s used in counting triangles. We extend $H_{\Delta}$ to the whole space smoothly and call the resulting Hamiltonian diffeomorphism $\phi_{H_{\Delta}}$. The reason for applying this Hamiltonian diffeomorphism is so that under the Lagrangian isotopy over $\delta$, the third side of the fiber triangle $u^{\prime\prime}$ is again in a linear Lagrangian